Absorption and Variable Costing Income Statements

During the first month of operations ended July 31, YoSan Inc. manufactured 11,900 flat panel televisions, of which 10,900 were sold. Operating data for the month are summarized as follows:

Sales $1,798,500
Manufacturing costs:
    Direct materials $928,200
    Direct labor 273,700
    Variable manufacturing cost 238,000
    Fixed manufacturing cost 119,000 1,558,900
Selling and administrative expenses:
    Variable $141,700
    Fixed 65,200 206,900

Required:

1. Prepare an income statement based on the absorption costing concept.

YoSan Inc.
Absorption Costing Income Statement
For the Month Ended July 31
Sales $
Cost of goods sold:
Cost of goods manufactured $
Inventory, July 31
Total cost of goods sold      
Gross profit $
Selling and administrative expenses
Income from operations $

Feedback

1. Sales - (cost of goods manufactured - ending inventory*) = Gross profit; gross profit - selling and administrative expenses = income from operations
*(Manufactured Units - Sold units) x (total manufacturing costs/manufactured units)

Learning Objective 1 and Learning Objective 2.

2. Prepare an income statement based on the variable costing concept.

YoSan Inc.
Variable Costing Income Statement
For the Month Ended July 31
Sales $
Variable cost of goods sold:
Variable cost of goods manufactured $
Inventory, July 31
Total variable cost of goods sold
Manufacturing margin $
Variable selling and administrative expenses
Contribution margin $
Fixed costs:
Fixed manufacturing costs $
Fixed selling and administrative expenses
Total fixed costs
Income from operations $.   
Answer #1

Answer 1:

Ending inventory = (Manufactured Units - Sold units) x (total manufacturing costs/manufactured units) = (11900 - 10900) * 1,558,900 / 11900

= $131,000

Income statement based on the absorption costing concept:

YoSan Inc. Absorption Costing Income Statement For the Month Ended July 31 Sales $1,798,500 1,00 $131,000 Inventory, July 31

Answer 2:

Variable cost of goods manufactured =  Direct materials + Direct labor +  Variable manufacturing cost

= 928200 + 273700 + 238000

= $1,439,900

Inventory, July 31 = (11900 - 10900) * 1439900 / 11900

= $121,000

Income statement based on the variable costing concept:

YoSan Inc Variable Costing Income Statement For the Month Ended July 31 Sales Variable cost of goods sold Variable cost of go